# Active Polymer Behavior in Two Dimensions: A Comparative Analysis of Tangential and Push-Pull Models

### v1

@@ IV Conclusions
At small bending rigidity, however, although the qualitative behavior is similar across all three models, the spiral transition in the tangential model is slightly shifted. This shift occurs because, either at high bending rigidity or in the spiral phase, the different definitions of the active force lead to similar structural and dynamical behavior by producing a tangential force along the filament. In contrast, at low bending rigidity, where the filaments are more flexible, the different definitions of activity influence the angle distribution between consecutive monomers, leading to shifts in the spiral and reentrant phase transitions.
Overall, we conclude that in two dimensions, when the mean active force is the same and the bending rigidity is sufficiently large, the tangential and push-pull models behave essentially identically. However, in three dimensions, we expect the tangential and push-pull models to exhibit different behaviors. In the three-dimensional case, increasing activity in the tangential model tends to lead to a collapsed state Bianco et al. (2018); Locatelli et al. (2021). As a result, the different definitions of activity in three dimensions are likely to produce distinct behaviors between the models.
@@
GJ and JPM contributed equally. GJ, JPM, JMR, CV, and DAMF designed the research. GJ and JPM conducted the research, establishing the methodology, performing the analysis, and validating the data. DAMF developed the software. CV and DAMF supervised the project. All authors contributed to writing and editing the manuscript, participated in discussions, and provided meaningful insights.
@@ Acknowledgements
D.M.F. and G.J. gratefully acknowledge support from the Comunidad de Madrid and the Complutense University of Madrid (Spain) through the Atracción de Talento program, grant 2022-T1/TIC-24007. C.V. acknowledges funding from MINECO under grants IHRC22/00002 and PID2022-140407NB-C21.

### v2

@@ IV Conclusions
At infinite dilution and small bending rigidity, the qualitative behavior is similar across all three models, with the only difference being a slight shift of the spiral transition in the T model. We have shown that, either at high bending rigidity or in the spiral phase, the different definitions of the active force lead to similar structural and dynamical behaviors. In contrast, at low bending rigidity, where the filaments are more flexible, the different definitions of activity have a stronger influence on the angle distribution between consecutive monomers, leading to shifts in the spiral phase. Overall, we conclude that in two dimensions, when the mean active force is the same and the bending rigidity is sufficiently large, the tangential and push-pull models behave essentially identically. However, in three dimensions, we expect the tangential and push-pull models to exhibit different behaviors. In the three-dimensional case, increasing activity in the tangential model tends to lead to a collapsed state Bianco2108 ; LocatelliPRL2021 , with respect to the push-pull model. As a result, the different definitions of activity in three dimensions are likely to produce distinct behaviors between the models.
V Supplementary materials
This section includes additional technical details that complement the findings discussed in the main manuscript. Specifically, it provides an in-depth explanation of the polymer models used, including the interaction potentials governing the behavior of bonded and nonbonded interactions in the polymer chains. These include both attractive and repulsive forces, as well as bending stiffness parameters. The document also includes further structural analysis, such as the calculation of the end-to-end distance and the gyration radius for different polymer configurations and models. In addition, supplementary figures illustrate the structural properties, such as the gyration radius and end-to-end distances, across various conditions, expanding on the results presented in the main text.
@@ V Supplementary materials
GJ and JPM contributed equally. GJ, JPM, JMR, CV, and DAMF designed the research. GJ and JPM conducted the research, establishing the methodology, performing the analysis, and validating the data. DAMF developed the software. EL and PM provided critical insights into the role of active forces in the spiral state, the behavior of long filaments, and its implications, as well as contributing to the overall revision of the manuscript. CV and DAMF supervised the project. All authors contributed to writing and editing the manuscript, participated in discussions, and provided meaningful insights.
@@ Acknowledgements
D.M.F. and G.J. gratefully acknowledge support from the Comunidad de Madrid and the Complutense University of Madrid (Spain) through the Atracción de Talento program, grant 2022-T1/TIC-24007. C.V. acknowledges funding from MINECO under grants IHRC22/00002 and PID2022-140407NB-C21. J.M.R. acknowledges financial support from the UCM predoctoral contract (call CT15/23).